It is now flu season at Patch-Em-Up emergency health services. Accordingly, the demand rate at the Patch-Em-Up clinic has increased to 50 patient arrivals per hour. In addition, the time that nurses spend with patients has increased, so Patch-Em-Up has made a number of changes. First, they hired more nurses, but they also assigned some of these nurses to perform an initial “triage” interview with patients before having the patient see a more highly trained nurse practitioner. Patch-Em-Up currently has 6 registered nurses working in the triage resource pool, plus 8 nurse practitioners (NPs). A triage nurse spends 5 minutes with each patient, on average, before sending the patient to the pool of NPs. An NP spends an average of 12 minutes with each patient.

1) Which resource pool is the bottleneck in this process? Select one:

a. Triage nurses

b. Nurse practitioners

c. Neither triage nor NP

2) What is the capacity of the process?

3) Given the data in the problem, what is the utilization of the NP resource pool?

4) Patch-Em-Up predicts that, by having a triage nurse spend one more minute with each patient, this can save two minutes per patient for the nurse practitioner, on average. If they make this change, will it change the utilization of the NP resource pool?

Select one:

a. Yes

b. No

c. It cannot be determined from the data provided
Answer #1

1) Nurse practitioners

As there are 6 triage nurses and each nurse takes 5 minutes with a patient, the total patients served by a triage nurses is 6*60/5 = 72 patients

As there are 8 NPs and each NP takes 12 minutes with a patient, the total patients served by NPs is 8*60/12 = 40 patients.

The bottleneck is the resource pool with the least capacity. As capacity of NPs is less, it is the bottleneck.

2) Capacity of the process is the capacity of the bottleneck. As we see from question 1, capacity of bottleneck which is Nurse Practitioners is 40 and hence the capacity of the process is 40 patients per hour

3) As NP resource pool is the bottleneck, the utilization of the same is 100%. We can also see that from Little's formula. As flow rate is 40 and capacity is 40, utilization is flow rate/capacity = 1 or 100%.

4) No.

If patients spend 1 more minute with each patient, total time spent by patient at triage nurse is 6 minutes. Hence total patients processed by triage nurses per hour is 6*60/6 = 60 patients.

Total time spent by each patient with NP decreases by 2 and becomes 10 minutes.  Hence total patients processed by NP is per hour is 8*60/10 = 48 patients.

As NP is still has the lowest capacity and is the bottleneck, utilization wont change for NP resource pool and will stay at 100%